1. Find the cost to run a 1200-watt microwave oven, if you make microwave popcorn in 5 minutes.

(price per kWh is 13.267 cents)




1
Expert's answer
2022-05-26T14:07:06-0400
"{\rm cost}={\rm price}\times {\rm energy\; used}"

"{\rm cost}={13.267\:{\rm (cents/kWh)}\times 1.2\:\rm kW\times \frac{5}{60}\: hr}\\
=1.33\:\rm cents"